OpenStudy (anonymous):

The area of a rectangle is x^2 - 2x - 15 and the length of the rectangle is x + 3. Find the width of the rectangle.

OpenStudy (across):

You're looking for a polynomial \(P(x)\) such that \(P(x)(x+3)=x^2-2x-15\).

OpenStudy (across):

Can you perform either long or synthetic division?

OpenStudy (anonymous):

how would i go about solving this im clueless

OpenStudy (anonymous):

yes

OpenStudy (across):

Try and simplify\[\frac{x^2-2x-15}{x+3}.\]

OpenStudy (anonymous):

x-5

OpenStudy (anonymous):

is the answer

OpenStudy (across):

That's right.